Two students are 320 meters apart and began walking toward one another at constant rates. One travels at 1.5 meters per second and the other at 1.7 meters per second. Ho long will they will walk until they meet?
Answer by mananth(16946)

Speed of A = 1.5m/s
speed of B = 1.7m/s
..
Since they are walking towards each other the combined speed = 1.5+1.7 = 3.2m/s
distance = 320m
...
Time taken = 320/3.2
100 seconds
...
m.ananth@hotmail.ca